graph each ellipse and find its foci
9X^2 + 4Y^2 =9
I am not sure how to do this since the 4 doesn't cancel out nicely.
Thank you,
Connie

Answer by lwsshak3(11628)

graph each ellipse and find its foci
9X^2 + 4Y^2 =9
I am not sure how to do this since the 4 doesn't cancel out nicely.
**
Standard form of equation for an ellipse with vertical major axis:
(x-h)^2/b^2+(y-k)^2/a^2=1, a>b, (h,k) being the (x,y) coordinates of the center.
For given equation: 9x^2+4y^2=9
divide both sides by 9
x^2+y^2/(9/4)=1
center: (0,0)
a^2=9/4
a=3/2
..
b^2=1
b=1
.
c^2=a^2-b^2=3/2-1=1/2
Foci: (0, 0±c)=(0,0±1/2)=(0,1/2) and (0,-1/2)
see graph below:
y=±(2.25-2.25x^2)^.5